# Animatable Gaussians: Learning Pose-dependent Gaussian Maps for High-fidelity Human Avatar Modeling

### v3

@@ Appendix D Additional Experiments
Quantitative Ablation Study on Parametric Template. We conduct a qualitative ablation study on the parametric template in Fig. 7 of the main paper. We also quantitatively compare the reconstructed parametric template with the naked SMPL-X model [59] on the animation accuracy in Tab. A. The numerical results are computed on the first 500 frames and the “22010708” camera view in the “avatarrex_lbn1” sequence from AvatarReX dataset [107].
Quantitative Ablation Study on Backbones. We additionally report quantitative results of StyleUNet [83], UNet [70] and MLPs in Tab. B to further prove the effectiveness of the introduction of 2D CNNs and 2D parameterization. The numerical results are computed on the 60-560 frames and the “Cam127” camera view in the “Actor02/Sequence1” from ActorsHQ dataset [28].
Quantitative Ablation Study on Pose Projection. We quantitatively ablate pose projection in Tab. C. It shows pose projection realizes more accurate animation under novel poses. The numerical results are computed on the 1109-1119, 1447, 1515 and 1578 frames where the testing poses are out of the distribution of training poses.
Number of Principal Components in Pose Projection. Fig. B shows the animation results with different numbers of principal components in the pose projection strategy. It demonstrates that although PCA can project a novel pose into the distribution of the training poses for better pose generalization as shown in Fig. 9 of the main paper, too few principal components may lose some fine-grained garment details. We empirically found that setting the number of principal components to 20 could produce both detailed and generalized animation.
Figure B: Ablation study on the component number in the pose projection.
View Number. We quantitatively and qualitatively show the animation results trained with 3 views, 6 views and 14 views in Tab. D and Fig. C. They demonstrate that our method also supports sparse-view input and can realize comparable high-fidelity results. The numerical results are evaluated on the first 500 frames and the “22070932” camera view of the “avatarrex_zzr” sequence from AvatarReX dataset [107].
Figure C: Animation results trained with different numbers of views.
Table D: Quantitative evaluation on different view numbers.
@@
Figure D: Failure cases. (a,c) Animation results by our method, (b,d) ground-truth images. Our method fails to model the motion of hairs.
Appendix E Failure Cases
Our method cannot model the physical motion of components that are not driven by the body joints, e.g., the hairs, as illustrated in Fig. D, since we model the whole body including clothes, hands and hairs as an entangled Gaussian representation. We leave for future work a disentangled and compositional representation for modeling the dynamics of different components of the character.

### v4

@@ IV-F2 Backbones
Figure 15: Ablation study on the component number in the pose projection.
IV-F3 Pose Projection
We evaluate the pose projection strategy by removing it, i.e., directly inputting the position map into the StyleUNet. Fig. 14 and Tab. VIII show the qualitative and quantitative animation results with and without the pose projection under novel poses, respectively. It demonstrates that direct extrapolation with the novel position map results in unreasonable 3D Gaussians, since no similar poses in the training dataset. In contrast, the pose projection guarantees that the reconstructed position maps (Eq. 17) lie within the distribution of training poses, leading to reasonable and vivid synthesized appearances.
IV-F4 Number of Principal Components in Pose Projection
Fig. 15 shows the animation results with different numbers of principal components in the pose projection strategy. It demonstrates that although PCA can project a novel pose into the distribution of the training poses for better pose generalization as shown in Fig. 14, too few principal components may lose some fine-grained garment details. We empirically found that setting the number of principal components to 20 could produce both detailed and generalized animation.
IV-F5 View Number
We quantitatively and qualitatively show the animation results trained with 3 views, 6 views and 14 views in Tab. IX and Fig. 16. They demonstrate that our method also supports sparse-view input and can realize comparable high-fidelity results.
Figure 16: Animation results trained with different numbers of views.
TABLE IX: Quantitative evaluation on different view numbers.
@@ IV-F5 View Number
V Discussion
V-1 Conclusion
We present Animatable Gaussians, a new avatar representation for creating lifelike relightable and animatable human avatars with highly dynamic, realistic and generalized appearances from multi-view RGB videos. Compared with implicit NeRF-based approaches, we introduce the explicit point-based representation, 3D Gaussian splatting, into the avatar modeling, and leverage powerful 2D CNNs for modeling higher-fidelity human appearances. Based on the proposed template-guided parameterization and pose projection strategy, our method can not only faithfully reconstruct detailed human appearances, but also generate realistic garment dynamics for novel pose synthesis. By introducing physically-based rendering into the avatar representation, our method can produce realistic avatar animation under different novel illuminations. Overall, our method outperforms other state-of-the-art avatar approaches, and we believe that the proposed 3D Gaussian splatting-based avatar representation will make progress towards effective and efficient 3D human representations.
Figure 17: Failure cases. (a,c) Animation results by our method, (b,d) ground-truth images. Our method fails to model the motion of hairs.
V-2 Limitation
Our method entangles the modeling of the human body and clothes, limiting to changing the clothes of the avatar for applications like virtual try-on. A possible solution is to separately represent the body and clothes with multi-layer 3D Gaussians as NeRF-based approaches [50, 125]. Moreover, our method relies on the multi-view input to reconstruct a parametric template, limiting the application for modeling loose clothes from a monocular video. Finally, Our method fails to model the physical motion of components that are not driven by the body joints, e.g., the hairs, as illustrated in Fig. 17, since we model the whole body including clothes, hands and hairs as an entangled Gaussian representation. We leave for future work a disentangled and compositional representation for modeling the dynamics of different components of the character.